OpenStudy (anonymous):

Help! giving out medals and will fan! A firecracker shoots up from a hill 145 feet high with an initial speed of 80 feet per second. Using the formula H(t) = -16t2 + vt + s, approximately how long will it take the firecracker to hit the ground? 13 seconds 11 seconds Nine seconds Six seconds

OpenStudy (anonymous):

six seconds

OpenStudy (anonymous):

Thanks! how'd you get that?

OpenStudy (anonymous):

0 = -16t^2+80t+145 Use the quadratic formula. t = 6.4 seconds

OpenStudy (anonymous):

was it correct

OpenStudy (anonymous):

yeah :) thanks!

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Function 1 is the graph Function 2 f(x) = -x^2 + 2x - 15 Function ____ has the larger maximum. (Put 1 or 2 in the blank space) Numerical Answers Expected

OpenStudy (anonymous):

im not sure but i would like a medal

OpenStudy (anonymous):

the maximum of f(x) is when its derivative d(f(x))/dx = 0 take the derivative of f(x). set it equal to 0 and solve for x. replace that value back into f(x) to get the maximum. look at the graph too see if that value is higher or lower

OpenStudy (anonymous):

So what would it be? I'm not really sure

OpenStudy (anonymous):

@JamesWolf

OpenStudy (anonymous):

f(x) = -x^2 + 2x - 15 derivative = -2x + 2 -15 maximum when this = 0 0 = -2x +2 -15 2x = -13 x(max) = -13/2 put this back into the original. and find the value of f(x) at the maximum x the graphs maximum looks to be around 1. is f(-13/2) bigger or smaller than 1?

OpenStudy (anonymous):

smaller?

OpenStudy (anonymous):

if thats what it is then thats what it is. idk I havnt put it back into the function f(x)